OpenStudy (anonymous):

Approximate S10 of the definite integral f(x)=sqrt(4-x^3) on [-1,1]

hero (hero):

Everything seems pretty straightforward with the exception of the S10 part

OpenStudy (anonymous):

ditto @Hero. S10?

OpenStudy (anonymous):

Your are suppose to use Simpsons rule to approximate the integral

OpenStudy (eyust707):

ahhh

OpenStudy (anonymous):

sorry for being unclear!

hero (hero):

I'm not a big fan of Simpson Rule

hero (hero):

I'm a fan of punching things into calculators

OpenStudy (anonymous):

i agree :) but they need us to "know the concept"

OpenStudy (anonymous):

It's just plugging in!! Seriously!

OpenStudy (anonymous):

You have the equation somewhere on hand?

OpenStudy (anonymous):

b = 1 a = -1

OpenStudy (anonymous):

ok just to clarify then the number of subintervals n is 10 ?

OpenStudy (eyust707):

yes I believe n = 10

OpenStudy (anonymous):

so subinterval length would be 1/5? (1-(-1))/10

hero (hero):

Oh S10 means no. of subintervals = 10

OpenStudy (eyust707):

\[\Delta x = {b - a \over n }\]

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

thanks! i was just hoping there was a simpler trick than the simpsons rule haha guess not

OpenStudy (anonymous):

\[1/3\left[ \sqrt(5) +4(\sqrt(4)+\sqrt(3) \right]\]

OpenStudy (eyust707):

\[\int\limits_{a}^{b} f(x) dx \approx {\Delta x \over 3} [f(x _{0}) + 4 f (x_{1} ) + 2f(x_{2})... + 4f(x_{n-1}) + f(x_{n})]\]

OpenStudy (anonymous):

nice formal definition @eyust707!...did you do that in this equation editor??

OpenStudy (eyust707):

Yea I looked at patrickJMT from youtube to get the formula and used the editor to type it up.

OpenStudy (anonymous):

For the life of me I can't figure out how to produce upright fractions in the editor.

OpenStudy (anonymous):

patrick is the man...wish he'd do more examples at times, but overall very solid teacher.

OpenStudy (eyust707):

For example try: "slope= {x over y}"

OpenStudy (anonymous):

over!!!! that's it?? lol

OpenStudy (eyust707):

yes sir

OpenStudy (anonymous):

nice. thanks man

OpenStudy (eyust707):

and the brackets tell it how much goes "over" lol

OpenStudy (eyust707):

although slope is y over x haha

OpenStudy (anonymous):

haha...got it : )

OpenStudy (eyust707):

oops

OpenStudy (anonymous):

\[i n \over my head\]

OpenStudy (eyust707):

hahahaaa

OpenStudy (eyust707):

try out "space"

OpenStudy (anonymous):

nice!!!!! all this time i was using the division button on my dang laptop!!

OpenStudy (anonymous):

\[\space\]

OpenStudy (anonymous):

lol...i'm going to use space in place of "smh" now!

OpenStudy (anonymous):

It's not overtly disrespectful.

OpenStudy (eyust707):

im confusedd?? lol

OpenStudy (anonymous):

\[\space\]

OpenStudy (eyust707):

\[I \space am \space confused \space \checkmark \]

OpenStudy (eyust707):

see the "space"

OpenStudy (eyust707):

lol

OpenStudy (anonymous):

See in response to your "I'm confused" I sent a "blank" response instead of "SMH"

OpenStudy (eyust707):

ohhhh hahaha

OpenStudy (anonymous):

It's very underhanded

OpenStudy (eyust707):

that other one is "checkmark"

OpenStudy (anonymous):

got it \[\checkmark\] Is there some sort of FAQ on here for all the symbols or should I just google search the code?

OpenStudy (eyust707):

ya know ive never seen one on here about that but i bet somewhere on the net theres one

OpenStudy (eyust707):

the equation editor automatically puts the \'s and [] for you so it makes it easy

OpenStudy (eyust707):

check out Latex symbols

OpenStudy (eyust707):

on google

OpenStudy (anonymous):

Say what?

OpenStudy (anonymous):

latex?

OpenStudy (eyust707):

thats the code language the equation editor uses

OpenStudy (eyust707):

\[\huge \clubsuit \]

OpenStudy (eyust707):

"clubsuit" lol

OpenStudy (anonymous):

\[\bigodot\bigodot\] OHHHHHHHHHHH

OpenStudy (eyust707):

=D

OpenStudy (anonymous):

This is pretty neat. More so nerdy...but neat too! : ) I'd say I greatly appreciate your insight man!!

OpenStudy (eyust707):

I think they are more like: \[\huge \bigcup \bigcup \]

OpenStudy (anonymous):

Not my (dream) woman's.

OpenStudy (anonymous):

jk...

OpenStudy (anonymous):

Anyway, I'm getting off this thing for a while. Thanks again man! Take care

OpenStudy (eyust707):

you too